# A Comprehensive Benchmarking Analysis of Fault Recovery in Stream Processing Frameworks

### v1

@@ 6. Fault Recovery Characterization
This section shows results demonstrating how the injected failures impact the frameworks’ execution as well as how they react to the failures. We provide results on the application metrics and system resource utilization (see more in Section 5.3). This section intends to demonstrate the potential impacts of faults and their recovery on the different stream processing frameworks and their Quality of Service (QoS). In this section, we show first failures for the sake of visual clarity, Section 7 extends these characterizations providing results with up to 10 recurrent failures injected.
@@ 6.1. ALOS Recovery Characterization
Figure 3 shows how the executions with the different frameworks behave when pods executing worker instances are killed (represented by the red line). To have more comprehensive characterizations, we present results with three failures that occur periodically with an interval of 5 minutes. It is also important to note that the plots show different metrics (see y-axis labels) over time.
@@
When the failures occur, the worker instance impacted is lost. Then, the Kubernetes infrastructure creates a new pod that is added to the pool of worker instances. Moreover, the stream processing frameworks are expected to start scheduling computations to new worker instances. This is achieved in Flink and Kafka Streams by rebalancing the partitions to worker instances. In Spark Structured Streaming, the worker instances are added to the pool of executors and they start receiving tasks to compute.
Fault recovery observation 1: The frameworks recover correctly from the failures. From a quality of services perspective, it is noticeable that Flink executions recover quicker from failures in comparison to Kafka Streams and Spark Structured Streaming, which is evinced by a shorter downtime in the output throughput and quicker reduction of the latencies.
Performance observation 1: After failures, Flink can present fluctuations in the data consumed from Kafka. To the best of our knowledge, this phenomenon has not been previously reported. The variation in the amount of data ingested after failures may be attributed to the uncertainty of the timing of failures relative to the previously saved checkpoints. However, we have deferred further exploration of this phenomenon to future work, as it does not significantly impact the output throughput or the latency.
@@
Performance observation 2: Kafka Streams achieves a stable throughput with and without failures of one worker instance pod, but after failure has fluctuations in resources usage. Kafka Streams is the one that accumulates more lag after failures, which is mostly due to its default partition assignment strategy triggered after the failures. After the first failure, there are fluctuations in the CPU utilization and network traffic in the worker machines, which indicates a potential imbalance. This also relates to the partition assignment strategy where some worker instances in the consumer group are assigned to more partitions or partitions with more lag accumulated. See more in the resource consumption observation 2.
@@
Performance observation 3: Spark Structured Streaming’s dynamic definition of micro-batch intervals is finding suitable configurations. Dynamic micro-batching is evident in the first few minutes of the execution with the throughput and lag fluctuating significantly. Such fluctuations happen because the micro-batch interval is too large causing more data to be consumed and batches being triggered less frequently. After a few minutes, the micro-batch interval is self-regulated to close to 2 seconds, which is a configuration suitable to sustain the input throughput and maintain a reasonable latency (the p50 latency is comparable to the micro-batch interval). Still, Spark’s latency is significantly higher than the other frameworks that process the records based on events instead of micro-batches. This outcome is consistent with the related literature (van Dongen and van den Poel, 2021).
@@
Resource consumption observation 1: Spark Structured Streaming has the most dynamic CPUs utilization. This is due to its modern execution model that processes based on fine-grain tasks assigned to the spark’s executors (worker instances). Tasks are created within the so-called stages that execute the micro-batches, the number of such tasks is determined based on the number of partitions (consumed from Kafka and subsequent shuffle partitions between DAG’s stages). Such a dynamic execution model can be more scalable and suitable for effective load-balancing but at the cost of potential overheads of having to assign all the fine-grain tasks to the executors. It is also noticeable that, from the operating system view, Spark Structured Streaming execution is the one that allocates the highest memory space. [^5] This memory observation should not be taken as evidence because the way the Java virtual machine allocates memory and computes can vary across the frameworks. Therefore, the operating system view of the memory usage does not guarantee accuracy. Considering that memory consumption can have a heavy impact on performance and cost, this motivates future analysis to comprehend the root causes and potential mitigations. One direction is to measure the garbage collection (GC) effectiveness and performance, where the Spark’s GC was fine-tuned in related works (van Dongen and van den Poel, 2021). Moreover, this behavior could be a potential reason for Spark’s performance limits found in our previous work (Henning et al., 2024).
@@
Resource consumption observation 2: From a resource consumption perspective, Kafka Streams shows unstable behavior after the failures are injected. This is supported mostly by varied CPU utilization and network traffic in the worker nodes, which suggests some imbalance that we believe to be load-related when repartitioning after the failures. This outcome aligns with performance observation 2: the partition assignment strategy can be suboptimal after failures making some workers overloaded and others underutilized. Furthermore, from an operating system view, Kafka Streams execution is the one allocating least memory, which is a positive outcome from a resource efficiency view.
@@
Resource consumption observation 3: Flink shows consistent resource utilization under failures. There are variations caused by the failures, but the execution returns to a normal stable state. It is worth noting that Flink’s execution is the one causing more CPU utilization in the Kafka Broker servers, which are already large-sized instances. This may be attributed to Flink’s fine-grained event-based processing, which continuously generates outputs to the Kafka topic. We noted that Flink’s performance may rely more than other frameworks on the behavior of the Kafka brokers and expect to address this in a future analysis.
@@
Figure 4 shows the scenario where periodically two random pods executing worker instances are killed. We expect this scenario to be more challenging as more workers failing means that more internal state can be lost and more records need to be replayed from the input source and reprocessed. Compared to Fig. 3, Fig. 4 shows similar performance, fault recovery, and resource consumption observations. However, it is noticeable in Fig. 4 that as Kafka Streams’ execution accumulates more lag after the failures, its event latency increases dramatically.
@@
With a more critical failure as the one with two pods failing, the unbalances caused by Kafka Stream’s repartitions after failures evinced in Performance observation 2 and Resource consumption observation 2 become more critical and impactful. On the one hand, Kafka Streams strategy can evenly distribute the partitions to the available workers in some cases of repartitions after failures. On the other hand, there are many cases where the distribution of partitions is suboptimal. For instance, the results from Fig. 4 indicate that after the first failure, the repartition was suboptimal with some workers taking too long to process the lag accumulated. However, after the third failure, the repartition seems to be optimal in such a way that the lag is fast processed and the latency reduced. This is supported by the resource utilization data that shows that a worker node (green color) consumes more CPUs and has higher network traffic while the lag is being reduced. Potentially this specific node was running a pod of a worker instance assigned to a lagged partition that was able to process the lagged records. Therefore, Fault recovery observation 2 is that Kafka Streams’ fault recovery can have an inconsistent impact on the applications’ performance across recurrent failures.
Although four worker instances failure is a more critical scenario, the results are aligned with the insights from the failures of two worker instances. Therefore, for the sake of conciseness, we do not show here results with 4 pods being killed.
@@
Figure 5 shows the scenario where the failure injected is that random pods executing worker instances face an operating system crash (see Section 5.2 for further details and motivations for this scenario). As they become unavailable, the records being processed by them and their local state is lost. In contrast with the pods killed failure, in this case, no new pods are created. Therefore, the execution has to continue processing with fewer worker instances.
@@
Focusing on Fig. 4(a) after the failure is injected, in Fig. 4(b) significant drops in system utilization are noticeable in the machines executing the two crashed pods. On the other hand, the active worker instances in all three frameworks increase their resource utilization to cope with the application processing demand. When the failure injected is halted (the gray line), the executions can return to the normal state with crashed worked instances being re-added to the workers pool. From the performance and QoS standpoint evinced in Fig. 4(a), Fault recovery observation 3 is that the executions of the three frameworks recovered from a critical failure that reduces the resource availability, and when the failure is halted they returned to a comparable state as the one before the failure.
One can note that Flink’s execution is the one that recovers faster to a normal performance state. Kafka Streams is the one accumulating more lag and its execution took almost 3 minutes to normalize the latency. Spark Structured Streaming demonstrates fluctuations after the failure, which is due to reprocessing of the tasks that were assigned to the failed executors and due to the dynamic micro-batching. With fewer executors and consequently fewer worker instances resources, Spark Structured Streaming stabilized the micro-batching interval with a higher value of around 3 seconds. The increase in the micro-batching interval was seemingly necessary to cope with the input load speed. However, this results in a slightly higher latency. Moreover, it is noticeable that Spark Structured Streaming’s output throughput faces more fluctuations after the failure is injected.
After the failure is halted the executions with the three frameworks return to normal states. Comparing the different worker instances, Kafka Streams ends up with some noticeable imbalances in network traffic and CPU utilization (Fig. 4(b)).
@@ 6.2. EOS Recovery Characterization
Supporting exactly-once semantics (EOS) demands further configurations and deployments. One more Kafka broker was necessary for resource stability and the fault tolerance guarantee had to be configured in the frameworks. Moreover, as also done in the related literature (van Dongen and van den Poel, 2021; Wang et al., 2021) when testing EOS, the configuration of the isolation level is set to read only committed data, which prevents the reading of duplicated messages in downstream operators.
Considering that Spark Structured Streaming can not natively guarantee exactly-once when writing the output to a Kafka topic, we adopt the same strategy of the related literature (van Dongen and van den Poel, 2021) and do not present Spark Structured Streaming with exactly-once.
For conciseness, we show only the results with one pod kill that already provides the insights necessary to highlight Flink’s advantages and Kafka Streams’ shortcomings when EOS is needed. In Fig. 6 it becomes evident that Flink, compared to Kafka Streams, provides better performance and fault recovery, [^6] We do not show the results from machine system monitoring of resources because it becomes repetitive and without new insights w.r.t. the previous results. thus resulting in Fault recovery observation 4: Flink provides better performance and fault recovery for use cases that demand exactly-once guarantees.
Focusing on the output throughput, both frameworks’ executions are similar until the first failure occurs. Then, Flink recovers faster in less than a minute to a normal output throughput that becomes stable. On the other hand, Kafka Streams took more than one minute to reach a necessarily higher output throughput. Moreover, after the failure, Kafka Streams’ output throughput is unstable and subject to fluctuations. We believe that this is due to the lag accumulated and being too slowly processed. This aligns with Performance observation 2 and Fault recovery observation 2. However, the scenario with exactly once is a more critical one, which results in a stronger impact of Kafka Streams’ suboptimal implementations.
One can note that Flink achieved a much lower latency in Fig. 6. Kafka Streams stabilized in a latency above 2 seconds and after the first failure, the latency does not normalize due to the large amount of lagged records. After a failure, Kafka Streams’ latency exceeds the 10-second metric bucket limit (Henning et al., 2024), rendering it invisible in the y-range plotted. These observations in Kafka Streams are, to the best of our knowledge, novel findings w.r.t. related approaches covering Kafka Streams with EOS (van Dongen and van den Poel, 2021; Wang et al., 2021). This can be mostly due to the fact the Kafka Streams was only tested under failures in (van Dongen and van den Poel, 2021) with a single failure injected. Moreover, technical reasons for the higher latency are due to the records buffering in Kafka Streams’ repartition topic and that its partition assignment strategy has been changed since (van Dongen and van den Poel, 2021), which implies that the current strategy is suboptimal for rebalancing after failures.
7. Fault Recovery Stability
This section presents the results of fault recovery stability under recurrent failures injected. Note the different ranges and scales on the y-axis. For conciseness, we limit the discussion to the results with ALOS guarantees as those are the most complete and insightful.
@@ 7. Fault Recovery Stability
Although Kafka Streams’ failure recovery pattern is consistent across the different failures, it varies significantly for how long the patterns and their phases occur, and their impact on performance (e.g., lag peaks and throughput drops). Moreover, the fluctuations in lag and throughput have a detrimental impact on the latencies, evincing limitations in Kafka Streams’ fault recovery. On the one hand, in the recovery of the sixth failure shown in 6(b), the lag accumulated as well as the throughput are lower, and the latency increase is not disruptive with the p50 being kept within the boundaries. On the other hand, following the seventh failure, the performance degradation becomes critical. Consequently, the execution takes around 4 minutes to return to a state comparable to the one before the failure.
In terms of reasons for the observed Kafka Streams’ fault recovery instabilities, the current understanding is that it is mostly due to which worker instance the failure is injected (e.g., if the consumer group leader is affected or not) and how the subsequent(s) repartitioning(s) are enacted. On the one hand, this outcome is aligned with Fault recovery observation 2. On the other hand, we believe that this potentially problematic behavior of Kafka Streams should be further characterized and improved.
Figure 8 complements the results from Fig. 5. If we compare the recovery from the first failure shown in Fig. 5 to the five failures covered in Fig. 8, one can note that Flink and Spark Structured streaming recoveries are similar in all failures. Kafka Streams, on the contrary, result in failures having (once again) different impacts on its performance, mostly in terms of latency, lag accumulation, and lag reduction. We can then conclude that Fault recovery observation 2 also applies to worker instances crashes.
@@ 8. Fault Recovery Time
The precise fault recovery time is another relevant metric that complements the results from Section 6 and Section 7. Considering that there is a lack of approaches and tools to automate the measurements of fault recovery times across many failures, we designed a simple detector based on moving average and standard deviation to automatically detect the failure as well as the time to recovery. The detector extracts the expected average and standard deviation for both throughput and latency based on the regular processing behavior that follows the warm-up period. The time frame ranged between 2 and 5 minutes, and 3 and 6 minutes, for throughput and latency, respectively. Figure 9 exemplifies the trace of automatically detected failures (of 2-pods kill) and their recoveries.
In short, if the moving average and standard deviation exceeded a given threshold (equal for all frameworks), then the detector would mark it as the point of failure injection. Afterward, another point of failure was strictly marked every 5 minutes. Similarly, the recovery time was detected when the estimators returned within the same threshold for a larger time frame, since once stability is achieved, throughput and latency should remain constant. Considering that the remaining experiments have a high correlation, for conciseness they are not shown here. We will provide the additional results and the parameters used for the thresholds in supplemental material.
Figure 9. Characterization of the automatic fault recovery times detection in the ALOS scenario of 2-pods kill.
Figure 10 summarizes the results of the fault recovery times. Noteworthy, throughput and latency (p90) recovery time are highly correlated. Flink achieves the fastest overall recovery time for both metrics. An exception is latency on 1-pod kill latency where Spark has a slightly lower median recovery ($m_{flink}=69.5$ seconds, $m_{spark}=68.5$ seconds) but at the cost of higher variability due to the observed spark’s fluctuations in throughput and latency.
Kafka Streams’ recovery times are longer and with higher variability, consistently with the discussion presented in Section 6. There are cases where full recovery was not achieved in the 5 minutes before the next failure was triggered, which is why the median can be high in some Kafka Streams’ results. Fault recovery observation 6 is that Flink and Spark Structured Streaming provide faster fault recoveries, and Flink demonstrates the highest level of stability.
Figure 10. Recovery times.
9. Concluding Remarks
Fault tolerance of data-intensive applications is of paramount importance in industry deployments. In this paper, we contribute with a more comprehensive evaluation method and fault recovery analysis. Such an evaluation was enabled by ShuffleBench’s (Henning et al., 2024) flexible deployment, parameters, and extension to fault recovery metrics.
9.1. Key Observations and Takeaways
We concluded that much has changed compared to the literature on how stream processing frameworks recover from failures. This motivates the need for benchmarking studies. The following are the key insights from the extensive experimental analysis:
  • Flink, Kafka Streams, and Spark Structured Streaming are resilient to different types and degrees of failures. However, the frameworks have different impacts on the applications running.

### v2

@@ 6. Fault Recovery Characterization
This section demonstrates how the injected failures impact the frameworks’ executions. The results cover application metrics and resource utilization (see more in Section 5.3). This section intends to demonstrate the potential impact of faults, their recovery on the different stream processing frameworks, and their Quality of Service (QoS). For visual clarity, The figures and plots provided for visual inspection depict either (i) the first three failures injected in the case of pod kills and (ii) the first failure followed by the frameworks’ response in the case of pod crashes. Section 7 extends the characterization with results from up to 10 injected failures.
@@ 6.1. ALOS Recovery Characterization
Figure 3 shows executions with the different frameworks with pods executing worker instances being killed (represented by the red line). It is also important to note that the plots show different metrics (see y-axis labels) over time. When the failures occur, the worker instance affected is lost. Then, the Kubernetes creates a new pod, which is added to the pool of worker instances. Moreover, the stream processing frameworks start scheduling computations to the new worker instances. This is achieved in Flink and Kafka Streams by rebalancing the partitions among the worker instances. In Spark Structured Streaming, the new worker instances start receiving tasks to compute.
Fault recovery observation 1: The frameworks recover correctly from the failures. However, when comparing the frameworks, it is noticeable that Flink executions recover quicker from failures. This observation is evinced by a shorter downtime in the output throughput and a faster latency recovery. Performance observation 1 is that after failures, Flink occasionally shows fluctuations in the input throughput. To the best of our knowledge, this observation has not been previously reported. Such fluctuations in the input throughput after failures can be due to the variation of the failures instant relative to the last saved checkpoints. Considering that these variations do not significantly impact the output throughput or the latency, further analysis is left for future work.
@@
Performance observation 2: Kafka Streams achieves a stable throughput with and without failures of one worker instance pod, but Kafka Streams’ executions are subject to unstable latencies. The Kafka Streams’ partition assignment strategy triggers a rebalance after failures, causing its executions to accumulate more lag. Rebalancing significantly increases event latency and shows fluctuations in the worker machines’ CPU usage and network traffic, indicating a potential imbalance. In fact, Kafka Streams’ performance following failures is mostly impacted by its partition assignment strategy, which unevenly distributes partition to the consumer group’s worker instances. Some worker instances may be assigned to more partitions or partitions with more accumulated lag. Then, some workers become overloaded while others are underutilized. [^4] Further analysis and potential optimizations in our follow-up paper (Vogel et al., 2024a).
@@
Performance observation 3: Spark Structured Streaming’s dynamic micro-batch intervals stabilize after a few minutes on suitable configurations. Dynamic micro-batching is noticeable in the first few minutes of execution due to fluctuations in throughput and lag, caused by an overly large micro-batch interval that results in more data being consumed and less frequent batches being triggered. Spark’s self-regulated micro-batch interval stabilizes at around 2 seconds per batch after a few minutes of execution. This configuration sustains the input throughput while maintaining reasonable latency. In our experience, the measured p50 latency is comparable to the micro-batch interval. Still, Spark’s latency is significantly higher than the other frameworks that process the records based on events instead of micro-batches. This insight is consistent with the related literature (van Dongen and van den Poel, 2021).
@@
Resource consumption observation 1: Spark Structured Streaming has the most dynamic CPUs usage. This is due to its modern execution model based on fine-grain tasks assigned to Spark’s executors (worker instances). Tasks are created within the so-called stages that execute the micro-batches. The number of tasks is defined based on the number of partitions (consumed from Kafka and shuffle partitions between the DAG’s stages). Such a dynamic execution model can be more scalable and provide load balancing, but it comes at the cost of potential scheduling overheads due to assigning all the fine-grained tasks to executors. From the operating system view, Spark Structured Streaming execution is the one that allocates the largest memory space. [^5] Considering that how Java Virtual Machine allocates memory and computes can vary across the frameworks, this memory observation should not be taken as evidence because the operating system’s view of memory usage does not guarantee accuracy. Considering that memory consumption can heavily impact performance, this motivates future analysis to comprehend the root causes and potential mitigations. One direction is to measure the garbage collection (GC) effectiveness and performance, where Spark’s GC was fine-tuned in related work (van Dongen and van den Poel, 2021). Moreover, this behavior could be a potential reason for Spark’s performance limits found in our previous work (Henning et al., 2024).
@@
Resource consumption observation 2: From a resource consumption perspective, Kafka Streams shows unstable behavior after the failures are injected. This is supported mostly by varied CPU usage and network traffic in the worker nodes, which suggests some imbalances aligned with performance observation 2. Furthermore, Kafka Streams’ executions allocate the least memory, which is a positive outcome from a resource efficiency view.
@@
Resource consumption observation 3: Flink shows consistent resource utilization under failures. Although the failures cause fluctuations, the executions return to a normal stable state. It is worth noting that Flink’s execution is the one causing more CPU usage in the large instances of Kafka Broker servers. This can be due to Flink’s fine-grained event-based processing, which continuously generates outputs to the Kafka topic. We noted that Flink’s performance may rely more than other frameworks on the behavior of the Kafka brokers. We intend to explore this insight further in a future analysis.
@@
Figure 4 shows the scenario where two random pods are periodically killed. We expect this scenario of more workers failing to be more challenging as more internal state is lost and more records need to be reprocessed by being replayed from the input source. Compared to Fig. 3, Fig. 4 shows similar performance, fault recovery, and resource consumption observations. However, it is noticeable in Fig. 4 that Kafka Streams’ execution accumulates more lag after the failures and its event latency increases dramatically.
@@
Kafka Streams’ execution shows good performance after failures in some cases, e.g., after the third failure demonstrated in Fig. 4 where the rebalance enables the lag to be processed quickly and the latency reduced. Figure 4(b) shows that a Kafka Streams’ worker node (indicated in green) consumes more CPUs and has higher network traffic while the lag is being reduced. This indicates that a worker instance pod on this specific node was assigned to a lagged partition and processed the lagged records. On the other hand, there are many cases where the distribution of partitions is suboptimal. For example, Fig. 4 shows that after the first failure, some workers take too long to process the accumulated lag compromising the recovery. Therefore, Fault recovery observation 2 is that Kafka Streams fault recovery’s impact on application performance varies across recurrent failures.
@@
We do not show here the results from four pods being killed because they are aligned with the insights from the failures of two worker instances (Vogel et al., 2024b). Figure 5 shows the scenario of failure injected where random pods executing worker instances crash (see more in Section 5.2). When pods become unavailable, the local state is lost. The execution must continue processing with fewer worker instances, as new pods are not created when pods crash.
@@
Significant drops in system utilization are noticeable after the failure in the machines executing the two crashed pods Fig. 5(b). Moreover, in the three frameworks, the remaining active workers increased their resource utilization to cope with the failure. When the injected failure is halted (the gray line), the executions can return to the normal state with crashed worked instances being re-added to the worker pool. From the performance and QoS standpoint evinced in Fig. 5(a), Fault recovery observation 3 is that the executions of the three frameworks recovered from a critical failure with fewer resource available, they also returned to a comparable state as the one before the failure when the failure is halted.
One can note in Fig. 5(a) that Flink’s execution is the one that recovers faster. Kafka Streams is the one accumulating more lag and its execution took almost 3 minutes to reduce the latency. After the failure is halted, Kafka Streams ends up with some noticeable imbalances in network traffic and CPU usage (Fig. 5(b)).
Figure 5(a) shows Spark Structured Streaming with fluctuations after the failure, which is due to the reprocessing of the failed executors’ tasks and its dynamic micro-batching. The micro-batching interval was increased to a higher value of around 3 seconds, attempting to cope with the same data volume and fewer processing resources. However, this caused a slight increase in Spark Structured Streaming’s latency and fluctuations in the output throughput.
@@ 6.2. EOS Recovery Characterization
Supporting exactly-once semantics (EOS) requires additional deployment considerations. One more Kafka broker was needed for resource stability and the fault tolerance guarantee had to be enabled in the frameworks. Moreover, the isolation level was configured to read only committed data to prevent the reading of duplicated messages in downstream operators, consistent with practices in related literature (van Dongen and van den Poel, 2021; Wang et al., 2021). Considering that Spark Structured Streaming can not natively guarantee exactly-once when writing the output to a Kafka topic. This relates to the output commit problem solved in Flink using the two-phase commit protocol and in Kafka Streams with custom sinks for rollback-recovery (Carbone et al., 2017; Fragkoulis et al., 2024). We do not show results from Spark Structured Streaming on exactly-once experiments, consistent with practices in related literature (van Dongen and van den Poel, 2021).
We show only the results with one pod kill that already provides the insights necessary to highlight Flink’s advantages and Kafka Streams’ shortcomings when EOS is needed. In Fig. 6 it becomes evident that Flink provides better performance and fault recovery than Kafka Streams, [^6] System monitoring results are not included in this section for the sake of conciseness, as there are no new insights from a resources perspective. thus resulting in Fault recovery observation 4: Flink provides better performance and fault recovery for use cases that demand exactly-once guarantees.
@@
One can note that Flink achieved a much lower latency in Fig. 6. Kafka Streams stabilized in a latency above 2 seconds and the latency does not normalize after the failure due to the significant lag. After failures, Kafka Streams’ latency is invisible in the y-range as it exceeds the 10-second bucket limit (Henning et al., 2024). The higher latency is due to the records buffering in Kafka Streams’ repartitioning topic.
Both frameworks’ executions are similar in the throughput until the first failure occurs. After the failure, Flink recovered faster in less than a minute. Kafka Streams took longer to reach a higher output throughput that remained unstable, which is due to the lag accumulated and being slowly processed. This aligns with Performance observation 2 and Fault recovery observation 2. Given that the scenario requiring exactly-once semantics is more critical, Kafka Streams’ suboptimal rebalance strategy has a greater impact.
7. Fault Recovery Stability
This section presents the results of fault recovery stability under recurrent failures injected. For the sake of visual clarity, note the different ranges and scales on the y-axis. For conciseness, we limit the discussion to the results with ALOS guarantees as those are the most complete and insightful.
Although Kafka Streams’ failure recovery pattern is consistent across the different failures, it varies significantly for how long the patterns and their phases occur, and their impact on performance (e.g., lag peaks and throughput drops). Moreover, the fluctuations in lag and throughput have a detrimental impact on the latencies, evincing limitations in Kafka Streams’ fault recovery. On the one hand, in the recovery of the sixth failure shown in 7(b), the lag accumulated as well as the throughput are lower, and the latency increase is not disruptive with the p50 being kept within the boundaries. On the other hand, following the seventh failure, the performance degradation becomes critical. Consequently, the execution takes around 4 minutes to return to a state comparable to the one before the failure.
In terms of reasons for the observed Kafka Streams’ fault recovery instabilities, the current understanding is that it is mostly due to which worker instance the failure is injected (e.g., if the consumer group leader is affected or not) and how the subsequent rebalances are enacted. On the one hand, this outcome is aligned with Fault recovery observation 2. On the other hand, we believe that this potentially problematic behavior of Kafka Streams should be further characterized and improved.
Figure 8 complements the results from Fig. 5. If we compare the recovery from the first failure shown in Fig. 5 to the five failures covered in Fig. 8, one can note that Flink and Spark Structured streaming recoveries are similar in all failures. Kafka Streams, on the contrary, result in failures having again different impacts on its performance, mostly in terms of latency, lag accumulation, and lag reduction. We can then conclude that Fault recovery observation 2 also applies to worker instances crashes.
@@ 8. Fault Recovery Time
Estimating fault recovery time is another relevant metric that complements the results from Section 6 and Section 7. Considering that there is a lack of approaches and tools to automate the measurements of fault recovery times across multiple failures, we designed a simple detector based on moving average and standard deviation to automatically detect the failure as well as the time to recovery. The detector extracts the expected average and standard deviation for the throughput based on the normal processing behavior. We define the normal processing period as the interval between the execution warm-up and the injection of the first failure, which can vary depending on the time required for warm-up and the timing of the first failure. In our analysis, the normal processing period is between the second and fifth minute after the start of execution across all frameworks.
In short, the detector marks a point of failure injection when the moving average of a window of 10 data points and its standard deviation are two times above the one from the normal processing period. The standard deviation threshold and window size are parametric values defined for all frameworks. Afterward, another point of failure was strictly marked every 5 minutes, which is the interval between failures used in our experiments.
The recovery time was detected when the values returned within a threshold of 15% for a time frame of at least 40 seconds. We expect throughput to stabilize after the recovery time, which should be only detected when the execution returns to a normal state. It is important to note that our approach for detection failures and their recovery was configured to automate our measurements. Considering that it is parametric and customizable, we expect it to be also applicable to wider scenarios. [^7] The tooling and all the parameters used are available as supplementary material (Vogel et al., 2024b). Moreover, we believe that there are no ground truths about these parameter values used. In fact, we focus more on having a formal parametric method to automate the measurements that is less arbitrary and avoids biased analysis. The values detected can be an estimation of the recovery time that we expect to be ascertained with visual inspections in the metrics traces. Figure 9 exemplifies the trace of automatically detected failures (of 2-pods kill) and their recoveries.
Figure 9. Characterization of our approach to automate measurements of fault recovery times in the 2-pods kill ALOS scenario. The stable behavior, used as a reference for training, is represented by a black rectangle. Failures are highlighted in red, while recovery detection is highlighted in green.
Figure 10 summarizes the results of the fault recovery times related to the application output throughput. Noteworthy, Flink achieves the fastest overall recovery time for both metrics. Kafka Streams’ recovery times are longer and with higher variability, which aligns with the discussion provided in Sections 6 and 7.
In the Kafka Streams scenario of 4 pods killed, a complete recovery was not achieved in the 5-minute window in four out of ten injected failures. The same event happened with one Kafka Streams failure in the scenario of 1 pod being killed. In such cases, the fault recovery time is defined as the time interval between failures. Defining the fault recovery time for these cases as the failure injection interval can impact the estimation because a complete recovery could take longer. [^8] The recovery accuracy can be improved with longer intervals between failures, where our follow-up paper provides preliminary results from such a scenario (Vogel et al., 2024a). Despite such a scenario, our method captured Kafka Streams’ fault recovery as the slowest due to its limited fault recovery.
Generally, Fault recovery observation 6 is that Flink and Spark Structured Streaming provide faster fault recoveries, and Flink demonstrates the highest level of stability. Considering that fault recovery time is also relevant in terms of the latency metric, we are currently measuring it, and preliminary results are already available in our follow-up paper (Vogel et al., 2024a).
Figure 10. Output throughput recovery times by framework.
9. Results Discussion
Our comprehensive experimental method demonstrates that stream processing frameworks significantly impact application performance and fault recovery. Given that fault recovery significantly affects the quality of service, it is a crucial criterion when selecting a stream processing framework for production deployment.  Section 9.1 shows our key insights from the experiments and Section 9.2 discusses the main findings.
9.1. Observations Summary and Takeaways
The following are key insights from our experiments and the observations discussed in Sections 6, 7 and 8:
  • Flink, Kafka Streams, and Spark Structured Streaming executions are resilient to recover from failures of killed and crashed worker instances, even from failures affecting many worker instances. Yet, the impact of their recoveries on the applications’ performance varies in terms of the application’s performance metrics, recovery time, and how stable the recoveries are across recurrent failures.